# Explaining Tournament Solutions with Minimal Supports

### v3

@@ 4.2 Weighted Uncovered Set
Even if MSs for $\operatorname{wUC}$ follow a structure similar to the unweighted case, we now show that deciding whether there exists a MS smaller than a given size is NP-complete.
@@ Appendix A Proofs for Section 4
The inclusion in NP is clear. To prove NP-hardness, we proceed with a reduction from the NP-complete Set Cover problem (karp1972reducibility). Given a set of $p$ element $U=\{e_{1},e_{2},\dots,e_{n}\}$ and a set $S$ of $q$ subsets of $U$, $S=\{s_{1},s_{2},\dots,s_{m}\}$ such that $\bigcup_{s\in S}s=U$, and an integer $k$, the set cover problem looks at the existence of a subset $D\subseteq S$ of size at most $k$ such that $\bigcup_{s\in D}s=U$. With this notation, we now prove that there exists a set cover of size less than $k\in\mathbb{N}$ if and only if there exists an MS of size less than $p+q+k$ for a specific tournament.
@@
$c_{e_{1}}$ $c_{e_{n}}$ 2 1 if $e_{i}\in s_{j}$ then 1 else 2
$c_{s_{1}}$ $c_{s_{m}}$ 0 if $e_{j}\in s_{i}$ then 1 else 0 1
@@
We now show that if there exists a set cover of $(U,S)$ of size less than $k$ then there exists an MS of $G(U,S)$ of size less than $p+q+k$ Suppose there exists a set cover $D$ of $(U,S)$ of size less than $k$. Let $\mathcal{X}\subseteq G$ such that: $\forall s\in D$, $\mu_{\mathcal{X}}(w,s)=2$, $\forall s\in S\setminus{D}$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},s)=1$ and $\forall e\in U$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},e)=1$. $|\mathcal{X}|=2|D|+1(p+q-|D|)=p+q+k$. $\forall s\in D$, $\mu_{\mathcal{X}}(w,c_{s})=2\geq\left\lceil(2+1)/2\right\rceil$. $\forall s\in S\setminus{D}$, $\exists s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(c_{s^{\prime}},c_{s})+\mu_{\mathcal{X}}(w,c_{s^{\prime}})=1+2\geq 2+1$. $\forall e\in U$, since $D$ is a cover set of (U,S), $\exists s\in D$ such that $\mu_{\mathcal{X}}(c_{s},c_{e})+\mu_{\mathcal{X}}(w,c_{s})=1+2\geq 2+1$. Hence, $w\in\text{NW}_{\operatorname{wUC}}(\mathcal{X})$ and there exists an MS contained in $\mathcal{X}$, i.e., of size at most $p+q+k$.
We now show the converse, if there exists an MS of $G(U,S)$ of size less than $p+q+k$ then there exists a set cover of $(U,S)$ of size less than $k$. Suppose there exists an MS $\mathcal{X}$ for $w\in\operatorname{wUC}(G)$ such that $|\mathcal{X}|=p+q+k$. Let $D=\{c:c\in\mathcal{C}\setminus{\{w\}},\mu(w,c)\geq\left\lceil(2+1)/2\right\rceil\}$ be the set of dominion of $w$. By minimality of an MS, $\forall c\in D$, $\forall c^{\prime}\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(c^{\prime},c)=0$ and $\forall c\in\mathcal{C}\setminus{D}$, $\exists!c^{\prime}\in D$, $\mu_{\mathcal{X}}(c^{\prime},c)=1$ and $\forall c^{\prime\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}}(c^{\prime\prime},c)=0$. Hence, p+q+k=2|D|+1(p+q|D|)=p+q+|D|p+q+k=2|D|+1(p+q-|D|)=p+q+|D| and $|D|=k$. However, since $\mu_{X}\leq\mu$ and $\forall e\in U$, $\mu(w,c_{e})=0$, $\forall e\in U$, $\exists c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu_{X}(c,c_{e})=1$ and $\mu_{X}(w,c)=2$. Additionally, $\forall e\in U$, $\forall c\in\mathcal{C}$, $\mu(c,c_{e})\geq 1$ and $\mu(w,c)\geq 2$ if and only if $\exists s\in S$ such that $c=c_{s}$ and $e\in s$. Thus, $\forall e\in U$, $\exists s\in S$ such that $e\in s$ and $s\in D$, i.e., $D$ is a set cover of $(U,S)$ of size $k$. ∎
@@ 20    $\text{rmed}\leftarrow\text{rmed}+\text{iteRm}$
When $w$ is a Condorcet winner, we are in the second case of Theorem 15. Indeed, in that case, $\sigma_{w}=m-1$, $\min_{c\neq w}\mu(w,c)=1$ and we have m1>1(m1)1m-1>1(m-1)-1. Additionally, $\left\lceil 1(1-1/m\right\rceil=0$ and $\mathcal{X}=m-1$. The other case is clear. ∎

### v4

@@ 4.2 Weighted Uncovered Set
Even if MSs for $\operatorname{wUC}$ follow a structure similar to the unweighted case, we now show that deciding whether there exists an MS smaller than a given size is NP-complete.
@@ Appendix A Proofs for Section 4
The inclusion in NP is clear. To prove NP-hardness, we proceed with a reduction from the NP-complete Set Cover problem (karp1972reducibility). Given a set of $p$ element $U=\{e_{1},e_{2},\dots,e_{p}\}$ and a set $S$ of $q$ subsets of $U$, $S=\{s_{1},s_{2},\dots,s_{q}\}$ such that $\bigcup_{s\in S}s=U$, and an integer $k$, the set cover problem looks at the existence of a subset $D\subseteq S$ of size at most $k$ such that $\bigcup_{s\in D}s=U$. With this notation, we now prove that there exists a set cover of size less than $k\in\mathbb{N}$ if and only if there exists an MS of size less than $p+q+k$ for a specific tournament.
@@
$c_{e_{1}}$ $c_{e_{p}}$ 2 1 if $e_{i}\in s_{j}$ then 1 else 2
$c_{s_{1}}$ $c_{s_{q}}$ 0 if $e_{j}\in s_{i}$ then 1 else 0 1
@@
We now show that if there exists a set cover of $(U,S)$ of size less than $k$ then there exists an MS of $G(U,S)$ of size less than $p+q+k$ Suppose there exists a set cover $D$ of $(U,S)$ of size less than $k$. Let $\mathcal{X}\subseteq G$ such that: $\forall s\in D$, $\mu_{\mathcal{X}}(w,s)=2$, $\forall s\in S\setminus{D}$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},s)=1$ and $\forall e\in U$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},e)=1$. $|\mathcal{X}|=2|D|+1(p+q-|D|)=p+q+|D|\leq p+q+k$. $\forall s\in D$, $\mu_{\mathcal{X}}(w,c_{s})=2\geq\left\lceil(2+1)/2\right\rceil$. $\forall s\in S\setminus{D}$, $\exists s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(c_{s^{\prime}},c_{s})+\mu_{\mathcal{X}}(w,c_{s^{\prime}})=1+2\geq 2+1$. $\forall e\in U$, since $D$ is a cover set of (U,S), $\exists s\in D$ such that $\mu_{\mathcal{X}}(c_{s},c_{e})+\mu_{\mathcal{X}}(w,c_{s})=1+2\geq 2+1$. Hence, $w\in\text{NW}_{\operatorname{wUC}}(\mathcal{X})$ and there exists an MS contained in $\mathcal{X}$, i.e., of size at most $p+q+k$.
We now show the converse, if there exists an MS of $G(U,S)$ of size less than $p+q+k$ then there exists a set cover of $(U,S)$ of size less than $k$. Suppose there exists an MS $\mathcal{X}$ for $w\in\operatorname{wUC}(G)$ such that $|\mathcal{X}|\leq p+q+k$. Let $D=\{c:c\in\mathcal{C}\setminus{\{w\}},\mu(w,c)\geq\left\lceil(2+1)/2\right\rceil\}$ be the set of dominion of $w$. By minimality of an MS, $\forall c\in D$, $\forall c^{\prime}\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(c^{\prime},c)=0$ and $\forall c\in\mathcal{C}\setminus{D}$, $\exists!c^{\prime}\in D$, $\mu_{\mathcal{X}}(c^{\prime},c)=1$ and $\forall c^{\prime\prime}\in\mathcal{C}\setminus{\{c^{\prime}\}}$, $\mu_{\mathcal{X}}(c^{\prime\prime},c)=0$. Hence, $|\mathcal{X}|=2|D|+1(|\mathcal{C}|-1-|D|)=2|D|+1(p+q-|D|)=p+q+|D|$ and since $|\mathcal{X}|\leq p+q+k$, we have $|D|\leq k$.
Since $\mu_{X}\leq\mu$ and $\forall e\in U$, $\mu(w,c_{e})=0$, $\forall e\in U$, we have $\exists c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu_{X}(c,c_{e})=1$ and $\mu_{X}(w,c)=2$. However, $\forall e\in U$, $\forall c\in\mathcal{C}$, $\mu(c,c_{e})\geq 1$ and $\mu(w,c)\geq 2$ if and only if $\exists s\in S$ such that $c=c_{s}$ and $e\in s$. Thus, $\forall e\in U$, $\exists s\in S$ such that $e\in s$ and $s\in D$, i.e., $D$ is a set cover of $(U,S)$. Hence, there exists a set cover of $(U,S)$ of size less than $k$. ∎
@@ 20    $\text{rmed}\leftarrow\text{rmed}+\text{iteRm}$
When $w$ is a Condorcet winner, we are in the second case of Theorem 15. Indeed, in that case, $\sigma_{w}=m-1$, $\min_{c\neq w}\mu(w,c)=1$ and we have m1>1(m1)1m-1>1(m-1)-1. Additionally, $\left\lfloor 1(1-1/m)\right\rfloor=0$ and $|\mathcal{X}|=m-1$. The other case is clear. ∎